# Improved Scaling Laws in Linear Regression via Data Reuse

### v1

@@ Abstract
Neural scaling laws suggest that the test error of large language models trained online decreases polynomially as the model size and data size increase. However, such scaling can be unsustainable when running out of new data. In this work, we show that data reuse can improve existing scaling laws in linear regression. Specifically, we derive sharp test error bounds on $M$-dimensional linear models trained by multi-pass stochastic gradient descent (multi-pass SGD) on $N$ data with sketched features. Assuming that the data covariance has a power-law spectrum of degree $a$, and that the true parameter follows a prior with an aligned power-law spectrum of degree $b-a$ (with $a>b>1$), we show that multi-pass SGD achieves a test error of $\Theta(M^{1-b}+L^{(1-b)/a})$, where $L\lesssim N^{a/b}$ is the number of iterations. In the same setting, one-pass SGD only attains a test error of $\Theta(M^{1-b}+N^{(1-b)/a})$ (see, e.g., Lin et al., 2024). This suggests an improved scaling law via data reuse (i.e., choosing $L>N$) in data-constrained regimes. Numerical simulations are also provided to verify our theoretical findings.
@@ 1 Introduction
A line of recent work establishes provable scaling laws in simplified settings such as linear regression  (see, e.g., Lin et al., 2024; Paquette et al., 2024, other related works will be discussed later in Section 6). Specifically, they consider an infinite-dimensional linear regression problem, where an $M$-dimensional linear model is trained by one-pass stochastic gradient descent (SGD) on $N$ Gaussian-sketched samples. Under power-law assumptions on the spectra of the data covaraince and the prior covariance, they show power-law type scaling laws in linear regression. However, their results are limited to one-pass SGD, where each sample is used once. In particular, Lin et al. (2024) attributed the nice, power-law type scaling laws to the implicit regularization effect of one-pass SGD (see Section 1 therein). It is unclear if scaling laws apply to other training algorithms, particularly those involving multiple passes of the data. Indeed, the scaling laws that only apply to one-pass method are not sustainable in a data-constrained regime.
@@
We assume that the data covariance and the prior covariance exhibit aligned power-law spectra with exponents $a$ and $b-a$, respectively (see Assumption 1C and 1D) (Lin et al., 2024; Paquette et al., 2024). We prove that multi-pass SGD achieves an excess test error of order $\Theta(M^{1-b}+L^{(1-b)/a})$ when $a>b>1$ and the number of SGD iterations $L\lesssim N^{a/b}$. This improves over the $\Theta(M^{1-b}+N^{(1-b)/a})$ bound for one-pass SGD (Lin et al., 2024) when $L>N$. In particular, when choosing the optimal number of iterations $L\eqsim N^{a/b}$, multi-pass SGD achieves an excess test error of order $\Theta(N^{(1-b)/b})$, in contrast to $\Theta(N^{(1-b)/a})$ for one-pass SGD in the data-constrained regime where $N\ll M^{b}$. Our results thus suggest that, to a certain extent, reusing data can improve the test performance of linear models in data-constrained regimes.
@@ 2 Setup
for some distribution ${P}$ on $\mathbb{H}\times{\mathbb{R}}.$ Given samples of the form $(\mathbf{x},y)$, instead of fitting a $d$-dimensional linear model, we train an $M$-dimensional sketched linear model with $M\ll d$. Namely, we consider linear predictors with $M$ parameters, defined as
@@
where $\mathbf{v}\in\mathbb{R}^{M}$ are the trainable parameters, and $\mathbf{S}\in\mathbb{R}^{M\times d}$ is some fixed sketching matrix. In this work, we consider Gaussian sketching, where the entries of $\mathbf{S}$ are drawn independently from $\mathcal{N}(0,1/M)$. Given a set of $N$ i.i.d. samples $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ from ${P}$, we train $f_{\mathbf{v}}$ via multi-pass stochastic gradient descent (multi-pass SGD), that is,
[…]
@@
where the expectation is over $(\mathbf{x},y)$ from ${P}.$ As an important component of our analysis, we also consider the gradient descent (GD) iterates
[…]
@@ 3 Main results
Assume the following conditions on the data distribution ${P}$.
@@
    for $\text{for $i\neq j$},\ \ {\mathbb{E}}[\langle\mathbf{v}_{i},\mathbf{w}^{*}\rangle\langle\mathbf{v}_{j},\mathbf{w}^{*}\rangle]=0;\ \ \text{and for $i>0$},\ \ {\mathbb{E}}[\lambda_{i}\langle\mathbf{v}_{i},\mathbf{w}^{*}\rangle^{2}]\eqsim i^{-b},\ \text{ for some $b>1$}.$
@@
Suppose Assumption 1 holds. Consider an $M$-dimensional linear predictor trained by (multi-pass SGD) on $N$ samples. Recall the decomposition in (3). Assume the initial stepsize $\gamma_{0}=\min\{\gamma,1/[4\max_{i}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}]\}$ for some $\gamma\lesssim 1/\log N$ and the number of effective steps ${L_{\mathtt{eff}}}\lesssim N^{a}/\gamma$. Then with probability at least $1-e^{-\Omega(M)}$ over $\mathbf{S}$
@@
    where the expectation is over the randomness of $\mathbf{w}^{*}$, $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ and $(i_{t})_{t=1}^{L}$. Moreover, when $a>b-1$, $\mathsf{Bias}$ and $\mathsf{Var}$ satisfy
@@
See the proof of Theorem 3.1 in Section A.2.1. A few comments on Theorem 3.1 are in order.
@@
[…]
@@ 3.1 Scaling laws for GD
Let Assumption 1 hold and $a>b-1$. Consider an $M$-dimensional linear predictor trained by (GD) on $N$ samples with stepsizes $\gamma_{0}=\min\{\gamma,1/[4\operatorname{tr}(\mathbf{S}\mathbf{X}^{\top}\mathbf{X}\mathbf{S}^{\top}/N)]\}$ for some $\gamma\lesssim 1$. Suppose $\sigma^{2}\eqsim 1$ and ${L_{\mathtt{eff}}}\lesssim N^{a}/\gamma$. With probability at least $1-e^{-\Omega(M)}$ over $\mathbf{S}$, the expected risk of $\bm{\theta}_{L}$ satisfies
@@
See the proof of Corollary 3.2 in Section A.2.2. From Corollary 3.2, we see that the variance error of GD is dominated by the sum of the approximation error and the bias error (i.e. $\mathsf{Var}\lesssim\mathsf{Approx}+\mathsf{Bias}$) when ${L_{\mathtt{eff}}}\gamma\lesssim N^{a/b}$. To achieve the optimal expected test error, we may choose $\gamma\eqsim 1$ and the number of effective steps ${L_{\mathtt{eff}}}\eqsim\min\{N^{a/b},M^{a}\}/\gamma\lesssim N^{a}$. Under this choice, we have
@@ 3.2 Scaling laws for multi-pass SGD
with probability at least $1-e^{-\Omega(M)}$. In contrast, Theorem 4.2 in Lin et al. (2024) proved that one-pass SGD with $\mathbf{v}^{o}_{0}=\bm{0}$, $\mathbf{v}^{o}_{t}=\mathbf{v}^{o}_{t-1}-\gamma_{t}\mathbf{S}\mathbf{x}_{t}(\mathbf{x}_{t}^{\top}\mathbf{S}^{\top}\mathbf{v}^{o}_{t-1}-y_{t})$ for $t\in[N]$ satisfies
@@
with probability at least $1-e^{-\Omega(M)}$, where ${N_{\mathtt{eff}}}:=N/\log N$. Here, all hidden constants depend only on $(a,b)$.
Several remarks on Corollary 3.3 are listed below.
@@
When $a\geq b>1$, Corollary 3.3 shows that multi-pass SGD achieves an excess test error of order $\Theta(M^{1-b}+({L_{\mathtt{eff}}}\gamma)^{(1-b)/a})$ when the number of effective SGD steps ${L_{\mathtt{eff}}}\lesssim N^{a/b}$, while one-pass SGD achieves an excess test error of order $\Theta(M^{1-b}+({N_{\mathtt{eff}}}\gamma)^{(1-b)/a})$. Therefore, the reused data across multiple passes (epochs) can be viewed as fresh data when the number of passes is smaller than $N^{a/b-1}$. For example, when $L=kN$ for some constant $k>1$, the test error achieved by $k$-pass SGD matches that of one-pass SGD trained on $kN$ $i.i.d.$ samples despite the training data being reused—aligning with the empirical observations in Muennighoff et al. (2023).
@@
$1𝑏𝑏\mathbb{E}[\mathcal{R}_{M}(\mathbf{v}_{L})]-\sigma^{2}\eqsim M^{1-b}+N^{\color[rgb]{0,0,1}{(1-b)/b}},$
@@
$1𝑏𝑎\displaystyle\mathbb{E}[\mathcal{R}_{M}(\mathbf{v}^{o}_{N})]-\sigma^{2}\eqsim M^{1-b}+{N_{\mathtt{eff}}}^{\color[rgb]{1,0,0}{(1-b)/a}}.$
@@ 4 Experiments
Figure 1: Multi-pass SGD versus one-pass SGD. In (a)—(c), multi-pass SGD is ran for $L\eqsim N^{a/b}$ steps. (a), (b), (d): SGD with geometrically decaying stepsizes; (c): SGD with constant stepsizes. We use linear functions to fit the excess test error in log-log scale. The fitted exponents ($k$) are close to the theoretical predictions in Corollary 3.3. The errorbars denote the ±$\pm 1$ standard deviation of the expected excess test error over $100$ i.i.d. samples of $(\mathbf{S},\mathbf{w}^{*})$. Parameters: $\sigma^{2}=1$, $d=10000$, $(a,b)=(2,1.5)$. (a), (b), (d): $M=1000$; (c): $N=10^{5}$.
@@ 5 Proof Overview
We provide an overview of the proof of Theorem 3.1 in this section. A full proof can be found in Section A.2.1. Let $\mathbf{v}^{*}=(\mathbf{S}\mathbf{H}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{H}\mathbf{w}^{*}$ and adopt the shorthand
@@
The main technical difficulty of proving Theorem 3.1 lies in bounding the bias, variance, and fluctuation error terms. For bias and variance, due to the non-commutativity of the population covariance ${\bm{\Sigma}}$ and the empirical covariance ${\bm{\widehat{\Sigma}}}$, we apply a covariance replacement trick (Lemma E.1; see also Lemma 7 in Pillaud-Vivien et al. (2018)) to replace the population covariance with the empirical covariance in the expression of bias and variance. For the fluctuation error, we mainly follow the proof techniques of Lemma 5 in Pillaud-Vivien et al. (2018), while also developing a more precise characterization of the magnitude of the GD outputs based on a leave-one-out argument (Lemma D.3). Although several upper bounds on the fluctuation error have been established for infinite-dimensional linear models (Pillaud-Vivien et al., 2018; Zou et al., 2022), the interaction between the sketching matrix $\mathbf{S}$ and the samples $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ in our setup introduces additional technical challenges (Lin et al., 2024). Moreover, we derive a novel lower bound on the fluctuation error that matches the upper bound up to a logarithmic factor in certain regimes.

### v2

@@ Abstract
Neural scaling laws suggest that the test error of large language models trained online decreases polynomially as the model size and data size increase. However, such scaling can be unsustainable when running out of new data. In this work, we show that data reuse can improve existing scaling laws in linear regression. Specifically, we derive sharp test error bounds on $M$-dimensional linear models trained by multi-pass stochastic gradient descent (multi-pass SGD) on $N$ data with sketched features. Assuming that the data covariance has a power-law spectrum of degree $a$, and that the true parameter follows a prior with an aligned power-law spectrum of degree bab-a (with $a>b>1$), we show that multi-pass SGD achieves a test error of $\Theta(M^{1-b}+L^{(1-b)/a})$, where $L\lesssim N^{a/b}$ is the number of iterations. In the same setting, one-pass SGD only attains a test error of $\Theta(M^{1-b}+N^{(1-b)/a})$ (see, e.g., Lin et al., 2024). This suggests an improved scaling law via data reuse (i.e., choosing $L>N$) in data-constrained regimes. Numerical simulations are also provided to verify our theoretical findings.
@@ 1 Introduction
A line of recent work establishes provable scaling laws in simplified settings such as linear regression  (see, e.g., Lin et al., 2024; Paquette et al., 2024, other related works will be discussed later in Section˜6). Specifically, they consider an infinite-dimensional linear regression problem, where an $M$-dimensional linear model is trained by one-pass stochastic gradient descent (SGD) on $N$ Gaussian-sketched samples. Under power-law assumptions on the spectra of the data covaraince and the prior covariance, they show power-law type scaling laws in linear regression. However, their results are limited to one-pass SGD, where each sample is used once. In particular, Lin et al. (2024) attributed the nice, power-law type scaling laws to the implicit regularization effect of one-pass SGD (see Section 1 therein). It is unclear if scaling laws apply to other training algorithms, particularly those involving multiple passes of the data. Indeed, the scaling laws that only apply to one-pass method are not sustainable in a data-constrained regime.
@@
We assume that the data covariance and the prior covariance exhibit aligned power-law spectra with exponents $a$ and bab-a, respectively (see Assumption 1C and 1D) (Lin et al., 2024; Paquette et al., 2024). We prove that multi-pass SGD achieves an excess test error of order $\Theta(M^{1-b}+L^{(1-b)/a})$ when $a>b>1$ and the number of SGD iterations $L\lesssim N^{a/b}$. This improves over the $\Theta(M^{1-b}+N^{(1-b)/a})$ bound for one-pass SGD (Lin et al., 2024) when $L>N$. In particular, when choosing the optimal number of iterations $L\eqsim N^{a/b}$, multi-pass SGD achieves an excess test error of order $\Theta(N^{(1-b)/b})$, in contrast to $\Theta(N^{(1-b)/a})$ for one-pass SGD in the data-constrained regime where $N\ll M^{b}$. Our results thus suggest that, to a certain extent, reusing data can improve the test performance of linear models in data-constrained regimes.
@@ 2 Setup
for some distribution P{P} on $\mathbb{H}\times{\mathbb{R}}.$ Given samples of the form $(\mathbf{x},y)$, instead of fitting a $d$-dimensional linear model, we train an $M$-dimensional sketched linear model with $M\ll d$. Namely, we consider linear predictors with $M$ parameters, defined as
@@
where $\mathbf{v}\in\mathbb{R}^{M}$ are the trainable parameters, and $\mathbf{S}\in\mathbb{R}^{M\times d}$ is some fixed sketching matrix. In this work, we consider Gaussian sketching, where the entries of $\mathbf{S}$ are drawn independently from $\mathcal{N}(0,1/M)$. Given a set of $N$ i.i.d. samples $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ from P{P}, we train $f_{\mathbf{v}}$ via multi-pass stochastic gradient descent (multi-pass SGD), that is,
$\begin{aligned} \mathbf{v}_{t}&:=\mathbf{v}_{t-1}-{\gamma_{t}}\big(f_{\mathbf{v}_{t-1}}(\mathbf{x}_{i_{t}})-y_{i_{t}}\big)\nabla_{\mathbf{v}}f_{\mathbf{v}_{t-1}}(\mathbf{x}_{i_{t}})\\ &=\mathbf{v}_{t-1}-{\gamma_{t}}\mathbf{S}\mathbf{x}_{i_{t}}(\mathbf{x}_{i_{t}}^{\top}\mathbf{S}^{\top}\mathbf{v}_{t-1}-y_{i_{t}}),&&t=1,\dots,L,\end{aligned}$ (multi-pass SGD)
@@
where the expectation is over $(\mathbf{x},y)$ from P.{P}. As an important component of our analysis, we also consider the gradient descent (GD) iterates
$\begin{aligned} \bm{\theta}_{t}&:=\bm{\theta}_{t-1}-\frac{\gamma_{t}}{N}\sum_{i=1}^{N}\big(f_{\bm{\theta}_{t-1}}(\mathbf{x}_{i})-y_{i}\big)\nabla_{\mathbf{v}}f_{\bm{\theta}_{t-1}}(\mathbf{x}_{i})\\ &=\bm{\theta}_{t-1}-\frac{\gamma_{t}}{N}\mathbf{S}\mathbf{X}^{\top}(\mathbf{X}\mathbf{S}^{\top}\bm{\theta}_{t-1}-\mathbf{y}),&&t=1,\dots,L,\end{aligned}$ (GD)
@@ 3 Main results
Assume the following conditions on the data distribution P{P}.
@@
    $\text{for $i\neq j$},\ \ {\mathbb{E}}[\langle\mathbf{v}_{i},\mathbf{w}^{*}\rangle\langle\mathbf{v}_{j},\mathbf{w}^{*}\rangle]=0;\ \ \text{and for $i>0$},\ \ {\mathbb{E}}[\lambda_{i}\langle\mathbf{v}_{i},\mathbf{w}^{*}\rangle^{2}]\eqsim i^{-b},\ \text{ for some $b>1$}.$
@@
Suppose Assumption˜1 holds. Consider an $M$-dimensional linear predictor trained by (multi-pass SGD) on $N$ samples. Recall the decomposition in (3). Assume the initial stepsize $\gamma_{0}=\min\{\gamma,1/[4\max_{i}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}]\}$ for some $\gamma\lesssim 1/\log N$ and the number of effective steps ${L_{\mathtt{eff}}}\lesssim N^{a}/\gamma$. Then with probability at least $1-e^{-\Omega(M)}$ over $\mathbf{S}$
@@
    where the expectation is over the randomness of $\mathbf{w}^{*}$, $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ and $(i_{t})_{t=1}^{L}$. Moreover, when a>b1a>b-1, $\mathsf{Bias}$ and $\mathsf{Var}$ satisfy
@@
See the proof of Theorem 3.1 in Section˜A.2.1. A few comments on Theorem 3.1 are in order.
@@
The assumption $\gamma\lesssim 1/\log N$ ensures that the initial stepsize $\gamma_{0}=\gamma$ with high probability. However, to guarantee the convergence of GD iterates, it suffices to have $\gamma_{0}\approx\gamma\leq 1/\|\mathbf{S}\mathbf{X}^{\top}\mathbf{X}\mathbf{S}^{\top}/N\|_{2}\overset{(i)}{\eqsim}1$. [^1] Step (i) follows from e.g., Theorem 4 and 5 in Koltchinskii and Lounici (2017). The additional $\log N$ factor and the assumption $\gamma_{0}=\min\{\gamma,1/[4\max_{i}\|\mathbf{S}\mathbf{x}_{i}\|_{2}^{2}]\}$ are technical conditions needed for analyzing the fluctuation error term. We leave the problem of relaxing these assumptions to future work.
@@
Relaxation of Assumption 1.
The Gaussian design in Assumption 1A can be relaxed to a sub-Gaussian design when establishing the upper bounds for $\mathsf{Bias},\mathsf{Var},\mathsf{Approx}$ in Theorem 3.1 and the upper bounds in subsequent corollaries. Moreover, the exact alignment of the eigenvectors of the prior and data covariance in Assumption 1D can be relaxed. We refer to Appendix A.3 for more details.
@@ 3.1 Scaling laws for GD
Let Assumption 1 hold and a>b1a>b-1. Consider an $M$-dimensional linear predictor trained by (GD) on $N$ samples with stepsizes $\gamma_{0}=\min\{\gamma,1/[4\operatorname{tr}(\mathbf{S}\mathbf{X}^{\top}\mathbf{X}\mathbf{S}^{\top}/N)]\}$ for some $\gamma\lesssim 1$. Suppose $\sigma^{2}\eqsim 1$ and ${L_{\mathtt{eff}}}\lesssim N^{a}/\gamma$. With probability at least $1-e^{-\Omega(M)}$ over $\mathbf{S}$, the expected risk of $\bm{\theta}_{L}$ satisfies
@@
See the proof of Corollary 3.2 in Section˜A.2.2. From Corollary 3.2, we see that the variance error of GD is dominated by the sum of the approximation error and the bias error (i.e. $\mathsf{Var}\lesssim\mathsf{Approx}+\mathsf{Bias}$) when ${L_{\mathtt{eff}}}\gamma\lesssim N^{a/b}$. To achieve the optimal expected test error, we may choose $\gamma\eqsim 1$ and the number of effective steps ${L_{\mathtt{eff}}}\eqsim\min\{N^{a/b},M^{a}\}/\gamma\lesssim N^{a}$. Under this choice, we have
@@ 3.2 Scaling laws for multi-pass SGD
with probability at least $1-e^{-\Omega(M)}$. Here, all hidden constants depend only on $(a,b)$.
In contrast, Theorem 4.2 in Lin et al. (2024) proved that one-pass SGD with $\mathbf{v}^{o}_{0}=\bm{0}$, $\mathbf{v}^{o}_{t}=\mathbf{v}^{o}_{t-1}-\gamma_{t}\mathbf{S}\mathbf{x}_{t}(\mathbf{x}_{t}^{\top}\mathbf{S}^{\top}\mathbf{v}^{o}_{t-1}-y_{t})$ for $t\in[N]$ satisfies
@@
with probability at least $1-e^{-\Omega(M)}$, where ${N_{\mathtt{eff}}}:=N/\log N$. Several remarks on Corollary 3.3 are listed below.
@@
When $a\geq b>1$, Corollary 3.3 shows that multi-pass SGD achieves an excess test error of order $\Theta(M^{1-b}+({L_{\mathtt{eff}}}\gamma)^{(1-b)/a})$ when the number of effective SGD steps ${L_{\mathtt{eff}}}\lesssim N^{a/b}$, while one-pass SGD achieves an excess test error of order $\Theta(M^{1-b}+({N_{\mathtt{eff}}}\gamma)^{(1-b)/a})$. Therefore, the reused data across multiple passes (epochs) can be viewed as fresh data when the number of passes is smaller than $N^{a/b-1}$. For example, when $L=kN$ for some constant $k>1$, the test error achieved by $k$-pass SGD matches that of one-pass SGD trained on kNkN $i.i.d.$ samples despite the training data being reused—aligning with the empirical observations in Muennighoff et al. (2023).
@@
[…]
@@
[…]
@@ 4 Experiments
Figure 1: Multi-pass SGD versus one-pass SGD. In (a)—(c), multi-pass SGD is ran for $L\eqsim N^{a/b}$ steps. (a), (b), (d): SGD with geometrically decaying stepsizes; (c): SGD with constant stepsizes. We use linear functions to fit the excess test error in log-log scale. The fitted exponents ($k$) are close to the theoretical predictions in Corollary 3.3. The errorbars denote the $\pm 1$ standard deviation of the expected excess test error over 100100 i.i.d. samples of $(\mathbf{S},\mathbf{w}^{*})$. Parameters: $\sigma^{2}=1$, $d=10000$, $(a,b)=(2,1.5)$. (a), (b), (d): $M=1000$; (c): $N=10^{5}$.
@@ 5 Proof Overview
We provide an overview of the proof of Theorem 3.1 in this section. A full proof can be found in Section˜A.2.1. Let $\mathbf{v}^{*}=(\mathbf{S}\mathbf{H}\mathbf{S}^{\top})^{-1}\mathbf{S}\mathbf{H}\mathbf{w}^{*}$ and adopt the shorthand notations
@@
The main technical difficulty of proving Theorem 3.1 lies in bounding the bias, variance, and fluctuation error terms. For bias and variance upper bounds, due to the non-commutativity of the population covariance $\bm{\Sigma}$} and the empirical covariance $^{\bm{\widehat{\Sigma}}}$, we apply a covariance replacement trick (Lemma E.1; see also Lemma 7 in Pillaud-Vivien et al. (2018)) to replace the population covariance with the empirical covariance in the expressions of bias and variance, as well as concentration properties of sub-Gaussian covariance to simplify their expressions. For the lower bounds, we show that a specific function of the empirical covariance commutes with the population covariance in expectation, and apply Von Neumann’s trace inequality.
For the fluctuation error, we follow the standard practice as in Pillaud-Vivien et al. (2018) and Aguech et al. (2000) to express the difference between the multi-pass SGD and GD trajectories, $\mathbf{v}_{t}-{\bm{\theta}}_{t}$ as a stochastic process (Eq. 18). We then bound the fluctuation error ${\mathbb{E}}[\|{{\bm{\Sigma}}}^{1/2}(\mathbf{v}_{L}-{\bm{\theta}}_{L})\|^{2}]$ through controlling the accumulated error of the stochastic process using Lemma D.2 and D.3, which involves a novel leave-one-out argument to control the model parameters. Although several upper bounds on the fluctuation error have been established for infinite-dimensional linear models (Pillaud-Vivien et al., 2018; Zou et al., 2022), the interaction between the sketching matrix $\mathbf{S}$ and the samples $(\mathbf{x}_{i},y_{i})_{i=1}^{N}$ in our setup introduces additional technical challenges (Lin et al., 2024). Moreover, we derive a novel lower bound on the fluctuation error that matches the upper bound up to a logarithmic factor in certain regimes by carefully controlling the accumulated variance from random sampling (i.e. the accumulated variance induced by the random indices $(i_{t})_{t=1}^{L}$).